# Controlled Training Data Generation with Diffusion Models

### v1

@@ 7.6 Image Guidance using Textual Inversion
where $\mathcal{E}$ is the VAE [39] image encoder and $\epsilon_{\theta}$ is the denoising UNet model from the Stable Diffusion model [61], and $x_{j}$ is sampled randomly from the set of available target images.
@@
Figure 16: Results of using Textual Inversion image guidance. We use the same plot as in Fig. 3. We add two additional methods using Textual Inversion image guidance described in Sec. 7.6. TI GAP is the Guided Adversarial Prompts with TI image guidance instead of the CLIP guidance. TI GP uses only TI guidance to find prompts. The TI guidance works on par with the CLIP guidance when used with adversarial optimization (TI GAP). However, using only TI guidance (TI GP) results in worse performance than using only text guidance with prompts found by the ALIA method [18].
@@ 7.7 Additional Analysis
8 Depth Estimation
8.1 Depth training details
Adversarial Optimization. The adversarial optimization was done with AdamW [46], learning rate of $5.0\times 10^{-4}$, weight decay of $1.0\times 10^{-3}$, and batch size of 8. The token embeddings at the start of optimization are randomly sampled from $\mathcal{N}(\mu_{emb},\sigma_{emb})$ where $\mu_{emb}$ and $\sigma_{emb}$ is the mean and standard deviation of all embeddings in the vocabulary. We set the early stopping threshold to 0.08 for the UNet model and 1.0 for the DPT model. Note that these models were trained with different losses, $\ell_{1}$ for the former and Midas loss [20] for the later. Adversarial optimization is performed with the same loss as was used for training these models. One run takes approximately 30 mins on one A100. We perform a total of 32 runs, to get 32 Adversarial Prompts for the UNet model and 30 runs for the DPT model. As the DPT model was trained on Omnidata, which is a mix of 5 datasets, we have 6 runs for each dataset. Different number of placeholder tokens were also used for each run as suggested in Fig. 5(b) of the main paper. For the DPT model, we do 1, 8, 16 tokens runs for each dataset and also 3 runs with 32 tokens for each dataset. For the UNet model, 4 runs of 1, 8 and 16 tokens each and 16 runs of 32 tokens were used, to get a total of 32 prompts. We also use a reduced number of denoising steps during optimization i.e., 5, as we found it to be more stable.
@@ 8.1 Depth training details
Generation. Generation is performed with the DDIM [72] scheduler and 15 sampling steps. We generate 80k images for the UNet model and 60k images for the DPT model for fine-tuning.
Fine-tuning. For fine-tuning, we optimize the UNet model with AMSGrad [60] with a learning rate of $5.0\times 10^{-4}$, weight decay of $2.0\times 10^{-6}$ and batch size 128. For the DPT model, a learning rate of $1.0\times 10^{-5}$, weight decay of $2.0\times 10^{-6}$ and batch size 32.
8.2 Additional Quantitative Results
Performance of non-SD baselines. In Tab. 6, we show the results for depth estimation for two additional baselines, deep augmentation [31] and style augmentation [23] that do not make use of generative models. Deep augmentation distorts a given image by passing it through an image-to-image model e.g., VAE [39], while perturbing its representations. Style augmentation involves involves applying style transfer to the original training images. They perform comparably to Adversarial Prompts.
@@ 8.2 Additional Quantitative Results
Table 6: Quantitative results on depth estimation. $\ell_{1}$ errors on the depth prediction task for a pre-trained U-Net and DPT model. (Lower is better. UNet losses are multiplied by 100 and DPT losses by 10, for readability. Note that the two models were trained with different losses, thus their numbers are not comparable to each other.). We evaluate on distribution shifts from Common Corruptions [29] (CC), 3D Common Corruptions [37] (3DCC) and cross-datasets (CDS), Replica [73]. The results from CC and 3DCC are averaged over all distortions and severity levels on Taskonomy. Our method is able to generate training data that can improve results over the baselines on several distribution shifts. Generations with AP (SDEdit) gives better results than AP under distribution shifts. Thus, also conditioning on the original image seems to be helpful for these shifts. For the DPT model, the trends are similar, AP performs better than the baselines. Deep augmentation and style augmentation do not make use of a generative model for generating extra data. They perform comparably to AP.
8.3 Additional Qualitative Results
Generations from all adversarial prompts & comparison of generations from different models. We show the generations from all Adversarial Prompts from the UNet model, without SDEDit (Fig. 19), with SDEdit (Fig. 20), and multi-iteration (Fig. 21). Additionally, we provide the generations from two DPT models, allowing us to assess the difference the model feedback has on generations. The first DPT model was only trained on Omnidata (Fig. 22) and second was trained on Omnidata with augmentations from CC and 3DCC and with consistency constrains [79] (Fig. 23). The quantitative results in the paper were reported only on the former DPT model.
@@ 8.4 Additional Analysis
8.4.1 Running multiple iterations of adversarial optimization vs a single iteration
Here, we provide additional analysis for the multi-iteration experiments in Fig. 5(b) in the main paper. We optimize for 4 prompts in each iteration and noticed that if the number of placeholder tokens in a given prompt is kept fixed throughout the iterations, the optimization to find new Adversarial Prompts becomes more difficult. However, if we increase the number of tokens at each iteration e.g., $1$ token per prompt for 1st, $8$ per prompt for 2nd, etc, we are able to consistently find new Adversarial Prompts. Thus, we aim to investigate the generalization of a given model to different Adversarial Prompts, e.g., is a model more likely to generalize to Adversarial Prompts with the same number of tokens.
@@ 8.4.1 Running multiple iterations of adversarial optimization vs a single iteration
8.4.2 CLIP image vs text guidance.
@@
We observed some differences in the generations with text vs. image guidance (Fig. 18, top). Text Guided Prompts generates corruptions that are more realistic that image Guided Prompts. For example, fog gets denser further away from the camera or around the floor when text Guided Prompts are used for generations. For image Guided Prompts, as it was guided by the image samples from CC where the corruption is applied uniformly over the image, it learns to also apply a more uniform corruption over the image.
@@
8.4.3 Generalization of Adversarial Prompts to different models.
@@
Table 7: Performance of a model on generated images from Adversarial Prompts from another model. For the UNet model we report the $\ell_{1}$ loss (×$\times 100$ for readability) and the DPT model, the Midas loss (×$\times 10$ for readability). The Adversarial Prompts attained from performing adversarial optimization on a UNet model and evaluated on the same model result in a loss of 7.63. Generations from the DPT model evaluated on the UNet model result in a loss of 7.17. Thus, the adversarial prompts found for one model seems to also be adversarial for another. We also report the loss on the original images for comparison.

### v2

@@ B.6 Image Guidance using Textual Inversion
where $\mathcal{E}$ is the VAE Kingma & Welling (2022) image encoder and $\epsilon_{\theta}$ is the denoising UNet model from the Stable Diffusion model Rombach et al. (2022), and $x_{j}$ is sampled randomly from the set of available target images.
@@
Figure 16: Results of using Textual Inversion image guidance. We use the same plot as in Fig. 3. We add two additional methods using Textual Inversion image guidance described in Sec. B.6. TI GAP is the Guided Adversarial Prompts with TI image guidance instead of the CLIP guidance. TI GP uses only TI guidance to find prompts. The TI guidance works on par with the CLIP guidance when used with adversarial optimization (TI GAP). However, using only TI guidance (TI GP) results in worse performance than using only text guidance with prompts found by the ALIA method Dunlap et al. (2023).
@@ B.7 Additional Analysis
Appendix C Depth Estimation
C.1 Depth training details
Adversarial Optimization. The adversarial optimization was done with AdamW Loshchilov & Hutter (2019), learning rate of $5.0\times 10^{-4}$, weight decay of $1.0\times 10^{-3}$, and batch size of 8. The token embeddings at the start of optimization are randomly sampled from $\mathcal{N}(\mu_{emb},\sigma_{emb})$ where $\mu_{emb}$ and $\sigma_{emb}$ is the mean and standard deviation of all embeddings in the vocabulary. We set the early stopping threshold to 0.08 for the UNet model and 1.0 for the DPT model. Note that these models were trained with different losses, $\ell_{1}$ for the former and Midas loss Eftekhar et al. (2021b) for the later. Adversarial optimization is performed with the same loss as was used for training these models. One run takes about 15 mins using one 80GB A100, on our internal cluster. We perform a total of 32 runs, to get 32 Adversarial Prompts for the UNet model and 30 runs for the DPT model. As the DPT model was trained on Omnidata, which is a mix of 5 datasets, we have 6 runs for each dataset. Different number of placeholder tokens were also used for each run as suggested in Fig. 6 of the main paper. For the DPT model, we do 1, 8, 16 tokens runs for each dataset and also 3 runs with 32 tokens for each dataset. For the UNet model, 4 runs of 1, 8 and 16 tokens each and 16 runs of 32 tokens were used, to get a total of 32 prompts. We also use a reduced number of denoising steps during optimization i.e., 5, as we found it to be more stable.
@@ C.1 Depth training details
Random prompts. In our proposed method, we optimize for $n$ embedding vectors, resulting in a prompt, $c$. Thus, to match this setting, from a Gaussian distribution fitted on the embeddings from the vocabulary, we sample $n$ random embeddings to create a random prompt to be used in the data generation.
Generation. Generation is performed with the DDIM Song et al. (2020) scheduler and 15 sampling steps. We generate 80k images for the UNet model and 60k images for the DPT model for fine-tuning. For the GP runs with SDEdit, we used strength 0.6, for the GAP runs, strength 0.9. See
During optimization, we use only 5 denoising steps, as it is more stable.
Fine-tuning. For fine-tuning, we optimize the UNet model with AMSGrad Reddi et al. (2019) with a learning rate of $5.0\times 10^{-4}$, weight decay of $2.0\times 10^{-6}$ and batch size 128. For the DPT model, a learning rate of $1.0\times 10^{-5}$, weight decay of $2.0\times 10^{-6}$ and batch size 32.
C.2 Additional Quantitative Results
C.2.1 Depth Estimation
Performance of non-SD baselines. In Tab. 7, we show the results for depth estimation for two additional baselines, deep augmentation Hendrycks et al. (2021) and style augmentation Geirhos et al. (2018) that do not make use of generative models. Deep augmentation distorts a given image by passing it through an image-to-image model e.g., VAE Kingma & Welling (2022), while perturbing its representations. Style augmentation involves involves applying style transfer to the original training images. They perform comparably to Adversarial Prompts.
@@ C.2.1 Depth Estimation
Table 7: Additional quantitative results on depth estimation. $\ell_{1}$ errors on the depth prediction task for a pre-trained U-Net and DPT model. (Lower is better. UNet losses are multiplied by 100 and DPT losses by 10, for readability. Note that the two models were trained with different losses, thus their numbers are not comparable to each other.). We evaluate on distribution shifts from Common Corruptions Hendrycks & Dietterich (2019) (CC), 3D Common Corruptions Kar et al. (2022) (3DCC) and cross-datasets (CDS), Replica Straub et al. (2019). The results from CC and 3DCC are averaged over all distortions and severity levels on Taskonomy. Our method is able to generate training data that can improve results over the baselines on several distribution shifts. Generations with AP (SDEdit) gives better results than AP under distribution shifts. Thus, also conditioning on the original image seems to be helpful for these shifts. For the DPT model, the trends are similar, AP performs better than the baselines. Deep augmentation and style augmentation do not make use of a generative model for generating extra data. They perform comparably to AP.
C.2.2 Classification with iWildCam
We show experimental results for several domain generalization baselines on iWildCam. The results for the synthetic data generation methods are based on the low-data regime, i.e., where we generate 4X less data. Our proposed methods, both AP and GAP, outperform these domain generalizaton baselines.
Method ResNet50 GroupDRO (Sagawa et al., 2019) ADA (Volpi et al., 2018) SagNet (Nam et al., 2021) L2D (Wang et al., 2021) IRM (Arjovsky et al., 2019) CausIRL (Chevalley et al., 2022) Agnostic Prompts AP GP GAP Accuracy 67.5 70.8 61.0 64.5 77.4 55.4 69.6 69.8 79.3 71.2 81.2
Table 8: Additional quantitative results on iWildCam. We compare our methods against several domain generalization baselines. GP and GAP still outperforms.
C.2.3 Training with additional generated data
We experimented with training on more generated data for Waterbirds and Depth estimation. We found that the trends are dataset specific. For Waterbirds, GAP underperforms GP with 2X the amount of extra data (relative to the original training data, see  Fig. 3 for the results for up to 1X extra data). However, for depth estimation GAP outperforms GP (see  Fig. 5 for the results up to $10^{4}$ extra data).
Dataset Waterbirds (Acc.↑) Depth ($\ell_{1}$ Err.×$\times 100$) Extra data 1X 2X $10^{4}$ $10^{5}$ GP 84.6 89.2 3.79 3.7 GAP 84 87.2 3.47 3.3
Table 9: Training with additional data. We extend the results in Fig. 3 and Fig. 5 to show the performance of our method with more generated data.
C.3 Additional Qualitative Results
Generations from all adversarial prompts & comparison of generations from different models. We show the generations from all Adversarial Prompts from the UNet model, without SDEDit (Fig. 19), with SDEdit (Fig. 20), and multi-iteration (Fig. 21). Additionally, we provide the generations from two DPT models, allowing us to assess the difference the model feedback has on generations. The first DPT model was only trained on Omnidata (Fig. 22) and second was trained on Omnidata with augmentations from CC and 3DCC and with consistency constrains Zamir et al. (2020) (Fig. 23). The quantitative results in the paper were reported only on the former DPT model.
@@ C.4 Additional Analysis
C.4.1 Running multiple iterations of adversarial optimization vs a single iteration
Here, we provide additional analysis for the multi-iteration experiments in Fig. 6 in the main paper. We optimize for 4 prompts in each iteration and noticed that if the number of placeholder tokens in a given prompt is kept fixed throughout the iterations, the optimization to find new Adversarial Prompts becomes more difficult. However, if we increase the number of tokens at each iteration e.g., $1$ token per prompt for 1st, $8$ per prompt for 2nd, etc, we are able to consistently find new Adversarial Prompts. Thus, we aim to investigate the generalization of a given model to different Adversarial Prompts, e.g., is a model more likely to generalize to Adversarial Prompts with the same number of tokens.
@@ C.4.1 Running multiple iterations of adversarial optimization vs a single iteration
C.4.2 CLIP image vs text guidance.
@@
We observed some differences in the generations with text vs. image guidance (Fig. 18, top). Text Guided Prompts generates corruptions that are more realistic that image Guided Prompts. For example, fog gets denser further away from the camera or around the floor when text Guided Prompts are used for generations. For image Guided Prompts, as it was guided by the image samples from CC where the corruption is applied uniformly over the image, it learns to also apply a more uniform corruption over the image. We also noticed that the diffusion model was not able to generate certain shifts e.g., compression artifacts, we leave further analysis to future work.
@@
C.4.3 Generalization of Adversarial Prompts to different models.
@@
Table 10: Evaluation performance of a model on generated images from Adversarial Prompts from another model (without fine-tuning). For the UNet model we report the $\ell_{1}$ loss (×$\times 100$ for readability) and the DPT model, the Midas loss (×$\times 10$ for readability). The Adversarial Prompts attained from performing adversarial optimization on a UNet model and evaluated on the same model result in a loss of 7.63. Generations from the DPT model evaluated on the UNet model result in a loss of 7.17. Thus, the adversarial prompts found for one model seems to also be adversarial for another. We also report the loss on the original images for comparison.
@@
Appendix D Broader Impact
We presented a method to control generative models to produce training data for supervised learning models. While our method is not particularly poised for negative use, it should be noted that powerful generative models are a general tool and the method has the potential to be used in ways authors did not intent. In addition, the data they are trained on may incorporate various societal biases or contain samples gathered in different ways from the internet. Furthermore, we will make our code and models publicly available. Thus, allowing for transparent inspection and safeguarding.